Which property is illustrated by the following statement? (46.2 + 25.8) +29.4= (25.8 + 46.2) + 29.4
Evaluate: 2y2(x+y) when x=1 and y=5

Answer by user_dude2008(1862)

Commutative property of addition


2y^2(x+y)=2(5)^2(1+5)=2(25)(6)=300